Solve by factoring: 16x2 - 14x + 1 = -6x
I need help with this . Can someone please explain what to do for me using this example. I need to know the steps as well as how you got the answer. Thankyou for you time. Your are appreshated.

Answer by rapaljer(4671)

16x^2 -14x+1=-6x

This is a quadratic equation, so start by setting it equal to zero by adding +6x to each side:

16x^2 -14x+6x +1=-6x+6x
16x^2 -8x+1=0

This is perfect square trinomial, so it factors like this:
(4x_____)(4x_____)=0
(4x-1)(4x-1)=0
4x-1=0
4x=1
x=1/4

In this exercise, you may need additional help with solving quadratic equations. However, the underlying problem may be being able to factor this trinomial. Either way, I have pages on my website that may be helpful to you.

If you need additional help with "Factoring" or "Solving Quadratic Equations", check out my website by doing a "Bing" search for my last name "Rapalje". Look for "Rapalje Homepage" at the top of the search list, and choose the link on my homepage "Basic, Intermediate and College Algebra: One Step at a Time." Click on "Basic Algebra", and look for the appropriate sections in Chapter 2. My non-traditional explanation should be easier to understand than your textbook!

R^2

Dr. Robert J. Rapalje, Retired
Seminole Community College
Altamonte Springs Campus
Florida